OpenStudy (mendicant_bias):

(ODE) Trying to solve a basic Differential Equation, prompt posted below momentarily.

OpenStudy (mendicant_bias):

\[\text{Solve the differential equation.}\]\[\frac{x^2}{y^2}\frac{dy}{dx}-\frac{\sqrt{y^2+4}}{x+1}=0\]

OpenStudy (mendicant_bias):

@dan815 , should I treat this like a separable equation, probably? Going to take a shot at it like that.

OpenStudy (mendicant_bias):

\[\frac{x^2}{y^2}\frac{dy}{dx}=\frac{\sqrt{y^2+4}}{x+1}\]

OpenStudy (mendicant_bias):

\[\frac{x^2}{y^2}\frac{dy}{dx}(x+1)=\sqrt{y^2+4}\]

OpenStudy (dan815):

its separable isnt it

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}(x^2)(x+1)=\sqrt{y^2+4}y^2\]

OpenStudy (mendicant_bias):

(Yeah, heh, I'm just still shaky on this and am basically looking for validation to make absolutely sure I'm headed in the right direction, heh, I really don't want to mess up at the beginning)

OpenStudy (dan815):

its separable so

OpenStudy (mendicant_bias):

Alright, but what do I do about this dy/dx business where the differentials would be attached to the wrong side?

OpenStudy (mendicant_bias):

Yeah, I am separating it, right?

OpenStudy (dan815):

might have to do some u sub or parts on that sqrt one

OpenStudy (mendicant_bias):

But I mean, the differentials would be attached to the wrong side, I think I figured it out

OpenStudy (dan815):